help ASAP!

when 25.0 ml of 0.100 M Co(NO3)2 is mixed with 45.0 mL 0.100 M KOH(aq), show by calculation whether a precipitate of Co(OH)2(s) will for. Ksp of Co(OH)2 is 1.3 * 10^-15
Answer #1

Co(NO3)2 + 2KOH  \to Co(OH)2 + 2KNO3

mole ratio of Co(NO3)2 and KOH = 1:2

mmoles of Co(NO3)2 = volume × molarity

= 25×0.100 = 2.5

mmoles of KOH = 45×0.100 = 4.5

Now, moles of Co(NO3)2 reacted = ( mmoles of KOH /2) = (4.5/2) = 2.25 .

mmoles of Co(OH)2 formed = mmoles of Co(NO3)2 reacted = 2.25.

Total volume of solution = (25+45) = 70 mL

Concentration of Co(OH)2

= (mmoles of Co(OH)2/ total volume)

= (2.25/70)

= 0.032 M.

Ksp of Co(OH)2 = 1.3×10-15

Let solubility of Co(OH)2 = S ( mol/L or M)

Co(OH)2 \to Co2+ + 2OH-

then, Ksp = [Co2+] [OH-]2

Or, Ksp = S × (2S)2

So, Ksp = 4S3

Or, 1.3×10-15 = 4S3

Or, s3 = (1.3×10-15/4)

Or, s3 = 3.25×10-16

Or, S = (3.25×10-16)1/3

Or, S = 6.875×10-6 M

As, concentration of Co(OH)2 in the solution is greater than the solubility of Co(OH)2 , so Co(OH)2 will form precipitation.